OpenStudy (anonymous):

Find the missing temperature information in each situation. 49. The high temperature is 208C. The low temperature is - 158C. What temperature is halfway between the high and the low?

OpenStudy (agreene):

Start by finding the total difference. 208 - 158 = 50 so that manes there is a total of 50 degrees difference. To find our midpoint we divide the difference in 1/2 50 / 2 = 25 that means that we add 25 degrees to the lower point and we have our midpoint! 158 + 25 = 183 degrees. to check, from our midpoint we should be able to add our 1/2 distance again and get our upper bound. 183 + 25 = 208 it checks :)

OpenStudy (anonymous):

omg thanks :DD

OpenStudy (anonymous):

lol i messed up ont he question it was supposed to be the high is 20 c and the low is 15 c i dont know where the 8's came from